You plan to make equal annual donations over a 6-year period (beginning 1 year from now) to a college for supporting poor students per year for 21 years, with the first scholarship to be awarded after 1 year from now. The cost of tuition at the college is $4000 per year and is expected to stay at that amount for 4 years. After that time (i.e., year 5), the tuition is expected to increase by 8% per year. If the college can invest the money at a rate of 10% per year, what size must the donations be? [Ans. A = $11,988]
Answer #1

ANSWER:

year 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21
tution fees per year with 8% increase from 5th year 4000 4000 4000 4000 4320 4665.6 5038.848 5441.95584 5877.312307 6347.497292 6855.297075 7403.720841 7996.018508 8635.699989 9326.555988 10072.68047 10878.4949 11748.7745 12688.67646 13703.77057 14800.07222

Now we will find the present worth of this cash flow at 10% rate.

pw = cash flow in year 1(p/f,i,n) + cash flow in year 2(p/f,i,n) + cash flow in year 3(p/f,i,n) + cash flow in year 4(p/f,i,n) + cash flow in year 5(p/f,i,n) + cash flow in year 6(p/f,i,n) + cash flow in year 7(p/f,i,n) + cash flow in year 8(p/f,i,n) + cash flow in year 9(p/f,i,n) + cash flow in year 10(p/f,i,n) + cash flow in year 11(p/f,i,n) + cash flow in year 12(p/f,i,n) + cash flow in year 13(p/f,i,n) + cash flow in year 14(p/f,i,n) + cash flow in year 15(p/f,i,n) + cash flow in year 16(p/f,i,n) + cash flow in year 17(p/f,i,n) + cash flow in year 18(p/f,i,n) + cash flow in year 19(p/f,i,n) + cash flow in year 20(p/f,i,n) + cash flow in year 21(p/f,i,n)

pw = 4,000(p/f,10%,1) + 4,000(p/f,10%,2) + 4,000(p/f,10%,3) + 4,000(p/f,10%,4) + 4,320(p/f,10%,5) + 4,665.6(p/f,10%,6) + 5,038.84(p/f,10%,7) + 5,441.95(p/f,10%,8) + 5,877.31(p/f,10%,9) + 6,347.49(p/f,10%,10) + 6,855.29(p/f,10%,11) + 7,403.72(p/f,10%,12) + 7,996.01(p/f,10%,13) + 8,365.7(p/f,10%,14) + 9,326.55(p/f,10%,15) + 10,072.68(p/f,10%,16) + 10,878.49(p/f,10%,17) + 11,748.77(p/f,10%,18) + 12,688.68(p/f,10%,19) + 13,703.77(p/f,10%,20) + 14,800.07(p/f,10%,21)

pw = 4,000 * 0.9091 + 4,000 * 0.8264 + 4,000 * 0.7513 + 4,000 * 0.683 + 4,320 * 0.6209 + 4,665.6 * 0.5645 + 5,038.84 * 0.5132 + 5,441.95 * 0.4665 + 5,877.31 * 0.4241 + 6,347.49 * 0.3855 + 6,855.29 * 0.3505 + 7,403.72 * 0.3186 + 7,996.01 * 0.2897 + 8,365.7 * 0.2633 + 9,326.55 * 0.2394 + 10,072.68 * 0.2176 + 10,878.49 * 0.1978 + 11,748.77 * 0.1799 + 12,688.68 * 0.1635 + 13,703.77 * 0.1486 + 14,800.07 * 0.1351

pw = 3,636.36 + 3,305.79 + 3,005.26 + 2,372.05 + 2,682.38 + 2,633.61 + 2,585.73 + 2,538.71 + 2,492.55 + 2,447.23 + 2,402.74 + 2,359.05 + 2,316.16 + 2,274.05 + 2,232.70 + 2,192.11 + 2,152.25 + 2,113.12 + 2,074.70 + 2,036.98 + 1,999.94

pw = 52,213.49

now we will find the annual equivalent amount.

n = 6

annual equivalent amount = pw(a/p,i,n) = 52,213.49(a/p,10%,6) = 52,213.49 * 0.2296 = 11,988.6

so the annual equivalent amount is $11,988